# Machine Vision-Based Assessment of Fall Color Changes and its Relationship with Leaf Nitrogen Concentration

### v2

@@ 11                    $c_{t}$ = $c_{t}$+$c_{i}$
Algorithm 1 Clustering of Point Cloud
@@
Figure 7: A custom program to select the points belonging to multiple classes - Green, Yellow, and Trunk. This instance showed an interactive window where the user selected points belonging to the trunk. The features of the points - including a*,b*,R, G, B, and their eigenvector and eigenvalues were recorded into a spreadsheet for further analysis.
@@
An alternative approach utilized a gradient boost classifier, an ensemble learning technique that combines multiple weak learners (typically decision trees) to create a stronger, more robust model (Bartlett et al., 1998). It sequentially builds new models, focusing on correcting the errors made by the previous ones, ultimately producing a more accurate and powerful predictive model (Natekin and Knoll, 2013). In this study, the gradient boost classifier model was created using scikit-learn (Van der Walt et al., 2014). The dataset was divided into 80% training and 20% testing sub-sets. The model required tuning of different hyperparameters. A strategic approach to varying model hyperparameters was employed, with one hyperparameter held constant while varying the others to analyze their individual effects on model accuracy. The hyperparameter learning rate was varied between 0.1-1, max depth was varied between 1-5, and n estimator was varied between 100-1000 and their performance on training and test dataset was analyzed. The set of hyperparameters that achieved the highest accuracy on the test data was then selected.
@@
To assess the performance of the yellowness estimation techniques, 50 validation trees were randomly selected and deleafed between the second and fourth trellis wires (a section of the trees, Figure 9). The leaves removed from each tree were collected in a bag and then later segregated manually into separate groups (“Green” or “Yellow”) for each tree. A leaf was identified as ’Green’ if more than half of the leaf area was green, and ’Yellow’ if otherwise. The precise weights of the yellow and green leaf groups were measured using a highly accurate Mettler PC 4000 (Mettler Toledo, Switzerland) weighing scale. This value served as the ground truth to assess the models’ performance. The yellowness index value was computed for all validation trees using Equation 1. Both Algorithm 1 and Gradient Boosting were evaluated against the ground truth data, with the superior-performing model based on this evaluation selected for further study.
@@
Figure 9: The image of the tree before and after deleafing. The leaves from the tree between the second and fourth trellis wire from the bottom (shown in transparent red box) were deleafed and collected.
2.3 Relationship between Yellowness Index and Leaf Nitrogen Concentration
@@
To investigate if the yellowness index had any relationship with the foliar nitrogen concentration on the trees, the yellowness index values were regressed against the leaf nitrogen concentration. The leaf nitrogen concentration was obtained using the standard Kjeldahl’s method (Kirk, 1950) by collecting 50 fully grown mid-shoot leaves from each tree, at the end of the growing season (Late July in 2021 and early August in 2023). The leaf nitrogen concentration would show the average nitrogen status of each tree and has been linked with various factors like canopy growth and canopy color. There are different recommendations on a good nitrogen level for trees depending on the variety (Cheng and Schupp, 2004; Cheng, 2010; Sallato, 2017; UniversityofArizona, ) . The leaf nitrogen level of 2-2.4% was regarded as a good nitrogen level in this study for Jazz™ apple trees based on these recommendations.
A one-way ANOVA analysis was conducted to examine the yellowness values across various weeks, classifying trees into either 3 or 5 groups according to their leaf nitrogen concentrations. In the 3-class classification, trees were categorized as ’Low’($N<2\%$),’Good’($2\%<N<2.4\%$), and ’High’($N>2.4\%$) . Meanwhile, the 5-class classification included ’VeryLow’($N<1.7\%$), ’Low’($1.7\%<N<2\%$), ’Good’($2\%<N<2.4\%$), ’High’($2.4\%<N<2.6\%$), and ’VeryHigh’($N>2.6\%$) categories. The ANOVA analysis was followed up by a Tukey HSD posthoc analysis to analyze the significant difference between pairs of groups if ANOVA showed a significant difference.
@@
A one-way ANOVA was performed to examine differences in the yellowness index among trees. This analysis was conducted separately for each week of the season, comparing trees grouped by their leaf nitrogen levels to determine if trees with different nitrogen statuses exhibited significant variations in their yellowness index throughout senescence. The trees were classified into 5 groups according to their leaf nitrogen concentrations: ’VeryLow’($N<1.7\%$), ’Low’($1.7\%<N<2\%$), ’Good’($2\%<N<2.4\%$), ’High’($2.4\%<N<2.6\%$), and ’VeryHigh’($N>2.6\%$). The ANOVA analysis was followed up by a Tukey HSD posthoc analysis to analyze the significant difference between pairs of groups if ANOVA showed a significant difference.
@@ 3.1 Tree Segmentation and Clustering
The original colored point cloud was clustered into green and yellow groups to quantify the yellowness of each tree canopy in the test orchard. As discussed in the methods section, Algorithm 1 (K-means based) and Gradient Boost classifier were used to classify leaves into these two groups. For the Gradient Boost model, learning rate, max depth, and n estimator were varied to identify the best set of hyperparameters (Figure 11). It can be seen from the figure that the model started to overfit the training data on increasing the hyperparameters, as indicated by higher performance on the training set with reduced performance on the test set (eg, training accuracy went from 0.85 to nearly 1, while testing accuracy decreased as $max\_depth$ of the trees was varied from 1 to 3, Fig 11b). The gradient boost method with the best set of hyperparameters achieved an accuracy of 78% on the test set with a learning rate of 0.1, maximum depth of 1, and a number of estimators as 100.
@@
The Gradient Boost model with the optimal hyperparameter configuration was compared against Algorithm 1 to determine the superior method among two. The models were evaluated using the validation data, detailed in Section 2.2.2. Figure 11 qualitatively shows the performance of the two models within cropped canopy regions (between the second and fourth trellis wire). Quantitative assessment of the performance of these models was performed using the Yellowness Indices (calculated with the proposed method as well as the manual estimation in the field) of the validation trees (Figure 12). It was found that both models achieved comparable classification accuracy on the validation dataset, as assessed by the $R^{2}$ value (0.69 for Algorithm 1 and 0.72 for gradient boost model), indicating a good fit between the ground truth and calculated yellowness. However, Algorithm 1 exhibited significantly slower processing speed compared to the Gradient Boost method. On the same point cloud, Gradient Boost model was approximately six times faster (with average segmentation time of 0.39 sec per tree) compared to Algorithm 1 (2.41 sec per tree). Consequently, Gradient Boost was selected as the optimal model for subsequent point cloud segmentation of all trees from all weeks of data collection. Figure 13 demonstrates the results from the gradient boost segmentation model for a test tree during weeks 1, 3, and 5.
@@
Figure 12: Performance of two classification/clustering models on the validation dataset; a) Estimated vs ground truth yellowness indexachieved by Algorithm 1; b) Same achieved by Gradient Boost; and c) Processing time (per image) on thirty-five test trees across all five weeks of data collection in 2023.
@@ 3.2 Yellowness Index and Leaf Nitrogen Concentration
The segmented point clouds were used to calculate yellowness indexfor each tree over individual weeks of image collection using equation 1. Spatial maps of these yellowness indices were depicted in Figure 14, which shows the pattern and progression of yellowing of the trees during fall, 2021 (Figure 14 a) and fall, 2023 (Figure 14 b). The darker points in these maps represent greener trees and the lighter points represent yellower trees. A consistent trend of increasing yellowness indexover time is noticed in both years. However, it was noticed that the yellowing trends were different between two years. In 2021, trees showed early transition to yellow in week 2 (around October 22), whereas in 2023, this change occurred later, nearly after three weeks (after November 10). This difference could potentially be attributed to relatively higher temperature during the growing season in 2021 compared to 2023 (Figure 15, AgWeatherNet ) as shown by the cumulative growing degree days (GDDS). Growing degree days (GDDs) are calculated by averaging the day’s maximum and minimum temperatures, subtracting a crop-specific base temperature below which growth ceases, and summing these values over the season (Miller et al., 2001). This method provides a more precise measure of plant growth stages than calendar days, as it accounts for the variable influence of temperature on developmental rates, allowing for more accurate predictions of phenological events regardless of annual temperature fluctuations. Prior studies byDe la Haba et al. (2014) and Kim et al. (2020) have highlighted the impact of higher temperatures on senescence, indicating that elevated temperatures can induce early senescence in foliage. The timing of full bloom differed by almost three weeks between the years, with blooms on April 13 in 2021 and May 2 in 2023. This difference in bloom timing correlates with the observed variations in the timing of tree yellowing, further supporting the role of GDDs in influencing key phenological stages. In both years, yellowing occurred approximately 29 weeks after full bloom. Combining temperature data with bloom dates could provide a prediction tool to estimate monitoring window for fall color changes in apple trees, which is essential for the practical application of the yellowness indexin optimizing N application.
@@
Figure 14: Spatial variation of yellowness indexover the study period in the test orchard; a) 2021 and b) 2023 datasets. The color bar on the right indicates the yellowness level, with darker shade representing greener trees and lighter shade representing yellower trees.
Figure 15: Growing degree days during 2021 and 2023 (AgWeatherNet, ). The red line shows the cumulative growing degree days for 2021 the green line shows that for 2023. The black line shows the average cumulative growing degree days between 2013 and 2022.
To explore the potential of color transformation in estimating N status in apple trees, correlation between leaf nitrogen concentrations and yellowness was investigated. Figure 16 shows leaf nitrogen concentrations plotted against the yellowness indexfor all five weeks of data collection in a)2021; and b) 2023. The points in scatter plots were categorized into five groups based on leaf nitrogen concentrations, as outlined in Section 3.2. In general, higher nitrogen levels in trees corresponded to lower yellowness index in both years ($r$ values ranging from -0.28 to -0.43 in 2021 and 0.16 and -0.64 in 2023), and an extended retention of greenness in tree foliage with higher N content compared to those with lower N levels. This finding confirms with existing literature emphasizing that trees with higher nitrogen levels tend to maintain green foliage color longer (Shi et al., 2021; Wen et al., 2018; Li et al., 2022).
@@
Figure 16: yellowness indexover weeks during the fall season of (a) 2021 and (b) 2023 in trees shown with Nitrogen % in the trees. (Very low N ($N<1.7\%$), Low N ($1.7\%<N<2.0\%$), Good N ($2\%<N<2.4\%$), High N($2.4\%<N<2.6\%$), Very high N ($N>2.6\%)$. The dates on the plots show the date when the data was collected in 2021 and 2023.
Qualitative resemblance of leaf nitrogen concentration and yellowness index can also be seen in Figure 17 (a – week 4 of 2021; b – week 5 of 2023). In both plots, red ellipses represent areas with higher leaf N and black ellipses represent areas with lower leaf N. It can be noticed in both of these plots that the area with lower leaf Nitrogen (right plots - black ellipses) has higher yellowness indices in the left plots and vice versa. These findings indicate that the yellowness index of trees estimated at different weeks during senescence can be a strong indicator of the nitrogen status in the tree, suggesting that leaf nitrogen content plays a crucial role in determining tree yellowness, which can be accurately captured by the yellowness index metric.
@@
Figure 17: Yellowneness Index (left plots) variation with leaf nitrogen concentrations (right plots) in (a) week 4 of 2021, and (b)week 5 of 2023. The colorbars on the left and right show the values of yellowess index and leaf nitrogen concentrations. The red ellipses show the sample areas with higher nitrogen concentrations, and the black ellipses show areas with lower nitrogen concentrations.
@@
The results from the one-way ANOVA suggested that a significant difference between the group means existed in weeks 1, 3, and 4 of 2021 and weeks 2, 4, and 5 of 2023, all at a significance level of $p<0.05$. Subsequently, post hoc analysis via TukeyHSD tests at the same significance level was conducted (Table 1). Rows shown in the table indicate that there existed a significant difference in group means between values in Column Group 1 and Group 2 at the given week. Significant differences in yellowness index between trees with different nitrogen levels were most prominent during weeks 2 and 4 in 2023 and week 4 in 2021.
These findings underscored the significance of timing the yellowness estimation in distinguishing trees with different nitrogen levels. To establish a more robust method for predicting these significant periods of observation for each cropping season, it’s essential to consider various factors such as growing degree days, precipitation, and average temperature. By identifying the critical time points (in terms of days after full bloom or growing degree days) when significant differences in leaf color occur between trees at different nitrogen concentrations, the yellowness index measured at those specific times would be crucial for accurately assessing tree nitrogen status.
@@
The results from this study demonstrated that a machine vision system could be used to identify the critical differences in the yellowing patterns of the trees and could be used as a predictor of nitrogen levels in apple trees. The outcomes and methodologies of this study offer a means to quantify tree color and determine tree nitrogen levels, which are crucial factors in guiding fertilizer application to individual trees. However, the system relies on a machine vision system which is susceptible to variations in external lighting conditions. Nevertheless, recent advancements in machine vision sensors have addressed low-light conditions using innovative techniques such as high dynamic range (Zhang et al., 2020, 2021) and Adversarial Networks (Chen et al., 2018). In addition, deep learning methods have notably enhanced imaging system resilience to diverse lighting environments (Li et al., 2021a, b). The algorithm employed in this study relies on labeled data for initially training the model to differentiate yellow and green points on trees, necessitating manual inputs. It’s important to note that this algorithm might require further calibration between different years and across tree varieties for optimal performance.

### v3

@@ 20                     $c_{t}=c_{t}\cup c_{i}$
Algorithm 1 Clustering of Point Cloud based on $a*$, $b*$ values
@@
Figure 7: A custom program to select the points belonging to multiple classes - Green, Yellow, and Trunk. This instance showed an interactive window where the user selected points belonging to the trunk. The features of the points - including a*,b*,R, G, B, and their eigenvector and eigenvalues (in a neighborhood) were recorded into a spreadsheet for further analysis.
@@
An alternative approach utilized a gradient boost classifier, an ensemble learning technique that combines multiple weak learners (typically decision trees) to create a stronger, more robust model (Bartlett et al., 1998).. It sequentially builds new models, focusing on correcting the errors made by the previous ones, ultimately producing a more accurate and powerful predictive model (Natekin and Knoll, 2013). In this study, the gradient boost classifier model was created using scikit-learn (Van der Walt et al., 2014). The dataset was divided into 80% training and 20% testing sub-sets. The model required tuning of different hyperparameters. A strategic approach to varying model hyperparameters was employed, with one hyperparameter held constant while varying the others to analyze their individual effects on model accuracy. The hyperparameter learning rate was varied between 0.1-1, max depth was varied between 1-5, and n estimator was varied between 100-1000 and their performance on training and test dataset was analyzed. The set of hyperparameters that achieved the highest accuracy on the test data was then selected.
@@
To assess the performance of the yellowness estimation techniques, 50 validation trees were randomly selected and deleafed between the second and fourth trellis wires (a section of the trees, Figure 9). The leaves removed from each tree were collected in a bag and then later segregated manually into separate groups (“Green” or “Yellow”) for each tree. A leaf was identified as ’Green’ if more than half of the leaf area was green, and ’Yellow’ if otherwise. The precise weights of the yellow and green leaf groups were measured using a highly accurate Mettler PC 4000 (Mettler Toledo, Switzerland) weighing scale. This value served as the ground truth to assess the models’ performance. The yellowness index value was computed for all validation trees using Equation 1. Both Algorithm 1 (K-means) and Gradient Boosting were evaluated against the ground truth data, with the superior-performing model based on this evaluation selected for further study. A timing test was also performed to evaluate the computation time of each of the model. Both models were run on a System76 Oryx Pro (System76, Inc., Colorado, USA) equipped with 16GB RAM, an 8GB NVIDIA GeForce RTX 3070 GPU (NVIDIA Corp., California, USA), and an Intel i7-11800H processor (Intel Corp., California, USA) at 2.30 GHz with 16 cores.
@@
Figure 9: Images of a sample tree before and after deleafing. The leaves of 50 sample trees within a section between the second and fourth trellis wire were removed (shown in transparent red box) and collected to estimate the ground truth yellowness index.
2.3 Yellowness Index and Leaf Nitrogen Concentration
@@
To investigate if the yellowness index had any relationship with the foliar nitrogen concentration on the trees, the yellowness index values were regressed against the leaf nitrogen concentration. The leaf nitrogen concentration was obtained using the standard Kjeldahl’s method(Kirk, 1950) by collecting 50 fully grown mid-shoot leaves from each tree, at the end of the growing season (Late July in 2021 and early August in 2023). The leaf nitrogen concentration represents the average leaf nitrogen concentration of each tree and is associated with canopy growth and canopy color. The leaf nitrogen level of 2-2.4% was regarded as an adequate range for nitrogen concentration based on referenced values for apples(Cheng and Schupp, 2004; Cheng, 2010; Sallato, 2017; UniversityofArizona, ).
@@
A one-way ANOVA was performed to examine differences in the yellowness index among trees. This analysis was conducted separately for each week of the season, comparing trees grouped by their leaf nitrogen levels to determine if trees with different nitrogen statuses exhibited significant variations in their yellowness index throughout senescence. The trees were classified into 5 groups according to their leaf nitrogen concentrations: ’VeryLow’($N<1.7\%$), ’Low’($1.7\%<N<2\%$), ’Good’($2\%<N<2.4\%$), ’High’($2.4\%<N<2.6\%$), and ’VeryHigh’($N>2.6\%$). The ANOVA analysis was followed up by a Tukey HSD posthoc analysis to analyze the significant difference between pairs of groups if ANOVA showed a significant difference between those groups.
@@ 3.1 Tree Segmentation and Clustering
The original color point cloud was clustered into green and yellow groups to quantify the yellowness of each tree canopy in the test orchard. As discussed in the methods section, Algorithm 1 (K-means based) and the Gradient Boost classifier were used to classify leaves into these two groups. For the Gradient Boost model, learning rate, max depth, and n estimator were varied to identify the best set of hyperparameters (Figure 11). The model started to overfit the training data on increasing the hyperparameters, as indicated by higher performance on the training set with reduced performance on the test set (e.g., training accuracy went from 0.85 to nearly 1, while testing accuracy decreased as $max\_depth$ of the trees was varied from 1 to 3, Fig 10b). The gradient boost method with the best set of hyperparameters achieved an accuracy of 78% on the test set with a learning ratee of 0.1, maximum depth of 1, and number of estimators as 100.
@@
Using the validation data (detailed in Section 2.2.2), the Gradient Boost model, with the optimal hyperparameter configuration, was compared against Algorithm 1. Figure 11 qualitatively shows the performance of the two models within the cropped canopy regions (between the second and fourth trellis wire). Quantitative assessment of the performance of these models was performed using the Yellowness Indices (calculated with the proposed method as well as the manual estimation in the field) of the validation trees (Figure 12). Both models were similar in their classification accuracy on the validation dataset with $R^{2}$ of 0.69 for Algorithm 1 and 0.72 for gradient boost model. This suggests there is a good fit between the ground truth (Yellowness indices of the validation trees) and calculated yellowness. However, Algorithm 1 had significantly slower processing speeds compared to the Gradient Boost method. On the same point cloud, the Gradient Boost model was approximately six times faster (with average segmentation time of 0.39 sec per tree) compared to Algorithm 1 (2.41 sec per tree). Consequently, Gradient Boost was used for subsequent point cloud segmentations. Figure 13 demonstrates the results from the gradient boost segmentation model for a test tree during weeks 1, 3, and 5.
@@
Figure 12: Relationship between estimated and manually calculated yellowness index for two classification/clustering models on the validation dataset; estimated vs ground truth yellowness index achieved by a) Algorithm 1; b) Gradient Boost; and c) Processing time (per image) on thirty-five test trees across all five weeks of data collection in 2023.
@@ 3.2 Yellowness Index and Leaf Nitrogen Concentration
The segmented point clouds were used to calculate the yellowness index for each tree over individual weeks using equation 1. The pattern and progression of yellowing of trees is depicted in spatial maps of these yellowness indices during fall 2021 (Figure 14a) and fall, 2023 (Figure 14b). The darker points in these maps represent greener trees and the lighter points represent trees with more yellow leaves. In both years we recorded a consistent increase in yellowness index over time. However, the progression of yellowing was different between the two years. In 2021, trees began the transition to yellow in week 2 (around October 22), whereas in 2023, this change occurred later in the year after November 10. This difference may be due to relatively higher temperature during the growing season in 2021 compared to 2023 (AgWeatherNet). Comparing years in growing degree days (GDD) provides a more precise measure of plant growth stages than calendar days, as GDDs account for the variable influence of temperature on developmental rates, allowing for more accurate predictions of phenological events regardless of annual temperature fluctuations (Dass and Rai, 2013). Prior studies by De la De la Haba et al. (2014) and Kim et al. (2020) have demonstrated that elevated temperatures can induce early senescence in foliage. In our research site, the timing of full bloom in 2021 was almost three weeks earlier than 2023. This difference in bloom correlates with the observed variations in the timing of tree yellowing, and further highlights the role of temperature influencing key phenological stages. In both years, yellowing occurred approximately 29 weeks after full bloom.
@@
Figure 14: Spatial variation of yellowness index over the study period in the test orchard; a) 2021 and b) 2023 datasets. The color bar on the right indicates the yellowness level, with darker shade representing greener trees and lighter shade representing yellower trees..
The correlation between leaf nitrogen concentration and yellowness index was determined for each week in 2021 and 2023 (Figure 15). The trees were categorized into five groups based on their leaf nitrogen concentrations as outlined in Section 2.3. In general, tree nitrogen levels were poorly correlated with yellowness index in both years (r values ranging from -0.28 to -0.43 in 2021 and 0.16 and -0.64 in 2023). We also observed an extended retention of greenness in trees that had high N content compared to those with lower N levels. Previous studies have shown strong correlations between color indices and leaf nitrogen concentration in various plant species. For example, Lee et al. (2019) reported a correlation coefficient of 0.85-0.91 between the SPAD readings and leaf Nitrogen concentration in apples, while El-Azazy (2018) found an $R^{2}$ of 0.81 for citrus trees. Haider et al. (2021) observed an even higher $R^{2}$ of 0.91 across multiple crops, though their system required crop-specific calibration. Our study’s lower correlation might be related to two factors: variable lighting conditions in natural settings, and a limited range of samples in the lower nitrogen concentration region.
@@
Figure 15: Correlation between N concentration and yellowness indexover weeks during the fall season of (a) 2021 and (b) 2023 in trees shown with Nitrogen % in the trees. (Very low N ($N<1.7\%$), Low N ($1.7\%<N<2.0\%$), Good N ($2\%<N<2.4\%$), High N($2.4\%<N<2.6\%$), Very high N ($N>2.6\%)$. The dates on the plots show the date when the data were collected in 2021 and 2023.
The variability among trees in leaf nitrogen concentration and yellowness index is presented in Figure 16 (a – week 4 of 2021; b – week 5 of 2023). In both plots, red ellipses highlight areas of the field with higher leaf N and black ellipses represent areas with lower leaf N. In both plots the area with lower leaf Nitrogen has higher yellowness indices in the left plots and vice versa.
@@
Figure 16: yellowness index (left plots) variation with leaf nitrogen concentrations (right plots) in (a) week 4 of 2021, and (b)week 5 of 2023. The color bars on the left and right show the values of yellowness index and leaf nitrogen concentrations. The red ellipses show the sample areas with higher nitrogen concentrations, and the black ellipses show areas with lower nitrogen concentrations.
@@
The results from the one-way ANOVA suggested that a significant difference between the group means existed in weeks 1, 3, and 4 of 2021 and weeks 2, 4, and 5 of 2023, at a significance level of $p<0.05$. (Table 1). To identify which specific nitrogen level trees differed, a subsequent Tukey HSD test was performed. The columns labeled "Group 1" and "Group 2" show the pairs of trees with those nitrogen levels exhibited significant differences during each week. Each row in the table represents a significant difference in group means between the values in the "Group 1" and "Group 2" columns for the specified week. Significant differences in yellowness index between trees with different nitrogen levels were most prominent during weeks 2 and 4 in 2023 and week 4 in 2021.
@@
While this study suggested a generally lower correlation, there were some weeks in which the correlations were moderate ( 0.4-0.6), indicating that there may be specific times when trees with different nitrogen concentrations display different yellowness indices. Capturing data during these periods could be useful in identifying trees at different nitrogen levels more effectively. To establish the appropriate timing for accurate and robust yellowness index assessment, various factors might need to be considered, such as temperature, growing degree days, precipitation, among others. However, it should be noted that, overall, the correlation remained poor, which suggests that other factors—such as differences in vigor of the trees, irrigation, or the presence of other nutrients—could also have contributed to the variation in yellowness index. By identifying these critical time points (in terms of days after full bloom or growing degree days) when significant differences in leaf color occur between trees at different nitrogen concentrations, the yellowness index measured could potentially be used for accurately assessing tree nitrogen status.
The results from this study demonstrated that a machine vision system could be used to identify differences in the yellowing patterns of trees. The outcomes and methodologies of this study offer a means to quantify tree color that could be used to determine tree nitrogen levels, which are crucial factors in guiding fertilizer application to individual trees. However, the system relies on a machine vision system which is susceptible to variations in external lighting conditions. Nevertheless, recent advancements in machine vision sensors have addressed low-light conditions using innovative techniques such as high dynamic range (Zhang et al., 2020, 2021) and Adversarial Networks(Chen et al., 2018). In addition, deep learning methods have notably enhanced imaging system resilience to diverse lighting environments(Li et al., 2021a, b). The algorithm employed in this study relies on labeled data for initially training the model to differentiate yellow and green points on trees, necessitating manual inputs. It’s important to note that this algorithm might require further calibration between different years and across tree varieties for optimal performance.